OpenStudy (anonymous):

I'm just clarifying this. I don't go to school. I have no one to help me at all. So I'm asking these questions. 6 out of 10 7x-5y=31 4x-7y=-3 A. (8, 5) B. (3, -2) C. (8, -2) D. (3, 5)

OpenStudy (anonymous):

just plug in the points

OpenStudy (anonymous):

multiply th first eqn by 4 and the second eqn by 7 and subtract them to get x and substitute x in one eqn to find y

hero (hero):

Do what the first guy suggested

OpenStudy (mrnood):

USe the METHOD that @FancyUsername has kindly shown oyu 5 times already This is the same concept as the others - just different numbers

OpenStudy (anonymous):

Same old elimination method. as mathavraj mentioned 7x-5y=31 (*4) 4x-7y=-3 (*7) and we get 28x - 20y = 124 28x - 49y = -21 we substract them and get 29y = 145 y = 5 4x - 7*5 = -3 4x = 32 x = 8 Check: 7 * 8 - 5 * 5 = 31 56 - 25 = 31 31 = 31 x= 8 y = 5 And as MrNood asked no direct aswer so choose the letter by yourself :P

OpenStudy (anonymous):

are you glad @MrNood ?

OpenStudy (mrnood):

I am glad that you, @FancyUsername are so competent in solving these questions. I am really sorry that OP will not make any attempt to solve for herself It is genuinely not 'help' to give answers when the concept has been explained so clearly. Education is not about tick boxes and tests, it's about learning. OP has learnt nothing from your posts.

OpenStudy (anonymous):

You cannot FORCE someone to learn @MrNood .No one can. If someone wants to learn he does it. But sometimes in this imperfect educations systems there are suituations where you just need correct combination of tick boxes and test answers.